How many liters of a 10% alcohol solution should be mixed with 12 liters of a 20% alcohol solution to obtain a 14% alcohol solution?
Answer by scott8148(5891)

x(10%) + 12(20%) = (12 + x)(14%)

clearing percentages ___ 10x + 240 = 168 + 14x

72 = 4x ___ 18 = x